Cumina Stores is an all-equity-funded retailer looking for some guidance on dividend policy. Over the last 3 years, the company has seen its revenues and net income increase and has maintained a dividend payout ratio of 40% of net income:

Year 3 Years ago 2 Years ago Most recent year
Revenues $1,000 $1,200 $1,500
Net Income $100 $120 $150
Depreciation $50 $60 $75

Over the same three year period, Cumina Stores has seen its cash balance decline from $ 60 million to $ 20 million.

a. Based on the information provided, estimate how much the firm reinvested in long-term assets and working capital over the 3 - Year period.  

b. Assume now that revenues, net income and depreciation are expected to grow 10% a year for the next two years. Working capital is expected to remain 25% of total revenues over the period and capital expenditures will be 200% of depreciation each year. If the firm wants to maintain its dividend payout ratio at 40% and increase its cash balance back to $ 60 million, how much debt (as a percent of reinvestment) will the firm have to take on for this to be feasible?
Answer #1

a).

Formula -3 Total 1500 3700 150 Year (n) Revenue Net income Depreciation Dividend Retention 75 Dep D = 40%*NI Re = NI-D 370 18

Net reinvestment over 3 years = retention + decline in cash balance = 222 + (60-20) = 262 million

Gross reinvestment over 3 years = net reinvestment + total depreciation = 262 + 185 = 447 million

b).

Formula R NI Dep 200%*Dep 25%*(Rn - Rn-1) 40%*NI Year (n) Revenue Net income Depreciation Capex Change in WC Dividend -1 1500

FCFE over next 2 years = Total dividends + cash balance increase = 138.6 + 40 = 178.6 million

Net reinvestment over next 2 years - total net income - FCFE = 346.5 - 178.6 = 167.9 million

Total reinvestment over next 2 years = capex - depreciation + change in WC = 346.5 -173.25 + 78.25 = 252 million

Debt required = total reinvestment - net reinvestment = 252 - 167.9 = 84.1 million

Debt required as a %age of total reinvestment = 84.1/252 = 33.37%